One grade of oil has 0.50% of an additive, and a higher grade has 0.75% of the additive. How many liters of each must be used to have 1000L of a mixture with 0.65% of the additive?: One grade of oil has 0.50% of an additive, and a higher grade has 0.75% of the additive. How many liters of each must be used to have 1000L of a mixture with 0.65% of the additive?
Answer by checkley77(3654)

.75x+.50(1000-x)=.65*1000
.75x+500-.50x=650
.25x=650-500
.25x=150
x=150/.25
x=600 L of 65% additive.
1000-600=400 L of 50% additive.
Proof.
.75*600+.5*400=650
450+200=650
650=650